[figure: 9.7 cm]

JAHN-TELLER EFFECT IN EPR SPECTRUM OF METAL-ORGANIC COMPLEX Cu(en)SO_{4}.2H_{2}O

A. Anders^{a}, A. Kaplienko^{a}, O.Kravchyna^{a}, V.Bondarenko^{a} A. Feher^{b}, M. Orendach^{b}, A.Orendachova^{b}, M.Kajnakova^{b}

aInst. of Low Temp. Phys. & Eng., NASU, Pr.Lenina 47, 61103 Kharkiv, Ukraine bFaculty of Science, P.J.Safarik University, Park Angelinum 9, 04154 Kosice, Slovakia

The temperature dependence of resonance spectrum of single crystal Cu(en)SO_{4}.2H_{2}O have been investigated in temperature range 2,3 - 60 K at the frequency 73 GHz. At T>15 K the angular dependence of spectrum is described by spin-Hamiltonian of axial symmetry with g_{z}=2,291 and g_{x}=g_{y}=2,067. At T<15 K the following effects are observed: i) the splitting of resonance line in external field orientation along axis b; ii)the spectrum symmetry decreasing up to rhombic one; iii) the considerable broadening of resonance line with temperature decreasing. These effects are discussed in frame of model of two nonequivalent Jahn-Teller centers with mutually connected dynamical distorsion in cell. The behaviour of resonance line at low temperatures is determined by critical broadening in short-range magnetic order region.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : A.Anders

Address for correspondence : Institute for Low Temperature Physics and Engineering of the Academy of Sciences of Ukraine Pr. Lenin 47, 61103 Kharkiv Ukraine

Email address : aanders@ilt.kharkov.ua